HenryWhite445:

What is the best approximation for the perimeter of a semicircle with a diameter of 35 inches? Use 3.14 to approximate pi. 54.95 in. 89.95 in. 109.90 in. 144.90 in.

Vocaloid:

Perimeter of a semicircle = (1/2)(pi)(d) + d

jhonyy9:

perimeter of a circle = 2pi*r = 2pi*(d/2) = pi*d perimeter of a semicircle = (perimeter of a circle) / 2 = pi *d /2 @Vocaloid excuse me - why is this (+d) ?

Vocaloid:

I’m on my phone so I can’t draw it But if you draw a half circle you have to include the flat part of the circle which adds the extra d

Vocaloid:

1 attachment
jhonyy9:

ok ty so much - i understand it now - there is the perimeter halfed and the diameter ok ty so much